These are Fraction Expressions. Could you explain how to work them?
5x/4 + 1/2 = x - 1/2
The answer is -4

Answer by Nate(3500)

5x/4 + 1/2 = x - 1/2 ~ Multiply by 4 to rid of fractions
4*5x/4 + 4/2 = 4x - 4/2
5x + 2 = 4x - 2
x = -4